OpenStudy (anonymous):

6b^3-24b^2+24b=0 solve

OpenStudy (anonymous):

b=0 b=2 b=2

OpenStudy (anonymous):

is one of the 2's a negeative??

OpenStudy (anonymous):

is one of the 2's a negeative??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok

OpenStudy (gw2011):

Here is how the answer b=0 and b=2 is solved: 6b(b^2-4b+4)=0 6b(b-2)(b-2)=0 6b=0 b=0 b-2=0 b=2 Therefore, b=0 and b=2 and you can see that -2 is not a solution to the problem